### v1

@@ 5.4 Computational experiments on real-world datasets
6 Conclusion and future directionsa
@@ Conclusion
First, implementation of additional algorithms. In our experiments, we implemented only $\mathit{odt}_{\text{rec}}$ and $\mathit{odt}_{\text{vec}}$, while $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ were not considered. We note that $\mathit{odt}_{\text{depth}}$ is challenging to parallelize compared with $\mathit{odt}_{\text{size}}$ , $\mathit{odt}_{\text{kperms}}$ and shares features with $\mathit{odt}_{\text{vec}}$. Nevertheless, both $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ has their advantages: $\mathit{odt}_{\text{depth}}$ has advantages for constructing large trees efficiently using a depth-first search strategy. It would be interesting to explore whether $\mathit{odt}_{\text{depth}}$ could be as effective for hyperplane decision trees as it is for axis-parallel decision trees, as reported in Brița et al. (2025).
Second, improved heuristics. The current two heuristics become ineffective when $K$ and $D$ are large, providing only marginal improvements or even deteriorating accuracy. Randomly generating hyperplanes is ineffective because the probability of finding feasible nested combinations is low. Future work could focus on developing more powerful heuristics for constructing hyperplane decision trees in high-dimensional or large-$K$ settings. One possible approach is to combine depth-first search with coreset selection methods.
@@
Third, conflicts between the general position assumption and categorical data. The correctness of $\mathit{hodt}$ relies on the general position assumption for data points. However, this assumption is reasonable primarily for numerical classification tasks, and real-world datasets often violate it—for example, categorical datasets with discrete features. In such cases, axis-parallel decision tree algorithms may be more appropriate than hyperplane decision tree algorithm. For categorical data, rather than creating $D$ splits for each data point, the algorithms of Brița et al. (2025) and Mazumder et al. (2022) efficiently exploit the structure of categorical features by eliminating duplicate splitting rules from the candidate sets. However, it remains unclear how to leverage categorical information in the hyperplane or hypersurface setting. Investigating ways to incorporate this assumption into hypersurface decision tree algorithms could lead to more efficient methods for categorical datasets.
@@
Fourth, algorithms with mixed splitting rules. In Part I of this paper, we proposed a generator for mixed-splitting rules. In principle, this generator can be used as input for $\mathit{odt}$ to produce decision tree with mixed-splitting rules; however, due to space constraints, we did not perform experiments with mixed-splitting rules. Future work could explore the construction of potentially more flexible models using these rules.
@@
Finally, extensions to random forests. Two key results from Breiman (2001) suggest promising directions for constructing random forests with hypersurface decision trees.shows that adding more trees does not lead to overfitting but converges to a limiting generalization error, while Theorem 2.3 establishes that generalization performance improves with the strength of individual classifiers. These insights align naturally with our framework: even in our Python implementation, the $\mathit{sodt}$ program can efficiently solve $1\times 10^{6}$ feasible nested combinations in tens of seconds. Moreover, flexible tree models often achieve stronger performance than axis-parallel trees. Future research could explore constructing large random forests with hypersurface decision trees, potentially yielding ensembles that outperform those based on axis-parallel trees and offering a promising direction for further study.
@@ Appendix A Algorithms
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
  2. 2.
    Output: Array of $\left(K,G\right)$-nested-combinations
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  8. 8.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  9. 9.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  10. 10.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  11. 11.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  12. 12.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  13. 13.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  14. 14.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  15. 15.
    return $\mathit{ncss}\left[K\right]$
Algorithm 5 $\mathit{nestedCombs}$
@@ A.2 Incremental ancestry relation matrix generator
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
  2. 2.
    Output: Array of $\left(K,G\right)$-nested-combinations
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  8. 8.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  9. 9.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  10. 10.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  11. 11.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  12. 12.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  13. 13.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  14. 14.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  15. 15.
        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // $\mathit{update}_{\text{arMat}}$ update $\left(k,G\right)$-nested-combinations and its associated ancestry relation matrix, defined in Algorithm 1 for hyperplane spliting rules.
  16. 16.
    return $\mathit{ncss}\left[K\right]$
Algorithm 6 $\mathit{\mathit{\mathit{nestedCombsFA}}}\left(K,G,\mathit{xs}\right)$
A.3 Experiments for $K>3$ decision tree

### v2

@@ 5.4 Computational experiments on real-world datasets
6 Conclusion and future directions
@@ Conclusion
First, conflicts between the general position assumption and categorical data. The correctness of $\mathit{hodt}$ relies on the general position assumption for data points. However, this assumption is reasonable primarily for numerical classification tasks, and real-world datasets often violate it—for example, categorical datasets with discrete features. In such cases, axis-parallel decision tree algorithms may be more appropriate than hyperplane decision tree algorithm. For categorical data, rather than creating $D$ splits for each data point, the algorithms of Brița et al. (2025) and Mazumder et al. (2022) efficiently exploit the structure of categorical features by eliminating duplicate splitting rules from the candidate sets. However, it remains unclear how to leverage categorical information in the hyperplane or hypersurface setting. Investigating ways to incorporate this assumption into hypersurface decision tree algorithms could lead to more efficient methods for categorical datasets.
@@
Second, algorithms with mixed splitting rules. In Part I of this paper, we proposed a generator for mixed-splitting rules. In principle, this generator can be used as input for $\mathit{odt}$ to produce decision tree with mixed-splitting rules; however, due to space constraints, we did not perform experiments with mixed-splitting rules. Future work could explore the construction of potentially more flexible models using these rules.
@@
Third, extensions to random forests. Two key results from Breiman (2001) suggest promising directions for constructing random forests with hypersurface decision trees. Theorem 1.2 shows that adding more trees does not lead to overfitting but converges to a limiting generalization error, while Theorem 2.3 establishes that generalization performance improves with the strength of individual classifiers. These insights align naturally with our framework: even in our Python implementation, the $\mathit{sodt}$ program can efficiently solve $1\times 10^{6}$ feasible nested combinations in tens of seconds. Moreover, flexible tree models often achieve stronger performance than axis-parallel trees. Future research could explore constructing large random forests with hypersurface decision trees, potentially yielding ensembles that outperform those based on axis-parallel trees and offering a promising direction for further study.
@@ A.1 Nested-combination generator
Algorithm 5 Nested Combination Generator ($\mathit{nestedCombs}(K,G,\mathit{xs})$)
1:Input: $\mathit{xs}$: data list, length NN; $K$: outer combination size; $G$: inner-combination size
2:Output: Array of $(K,G)$-nested-combinations
[…]
6:for $n\in\mathit{range}(0,N)$ do
7:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
8:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
9:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
10:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
11:  end for
12:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n+1}{G}$
13:  for $i\in\mathit{range}(C_{1},C_{2})$ do
14:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
15:    $\mathit{ncss}[k]\leftarrow\mathit{map}(\cup[i],\mathit{ncss}[k-1])\cup\mathit{ncss}[k]$ $\triangleright$ Update nested combinations
16:   end for
17:  end for
18:end for
19:return $\mathit{ncss}[K]$
@@ A.2 Incremental ancestry relation matrix generator
Algorithm 6 Nested Combination Generator with Ancestry ($\mathit{nestedCombsFA}(K,G,\mathit{xs})$)
1:Input: $\mathit{xs}$: data list, length NN; $K$: outer combination size; $G$: inner-combination size
2:Output: Array of $(K,G)$-nested-combinations
[…]
6:for $n\in\mathit{range}(0,N)$ do
7:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
8:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
9:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
10:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
11:  end for
12:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n+1}{G}$
13:  for $i\in\mathit{range}(C_{1},C_{2})$ do
14:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
15:    $\mathit{ncss}[k]\leftarrow\mathit{map}(\cup[i],\mathit{ncss}[k-1])\cup\mathit{ncss}[k]$ $\triangleright$ Update nested combinations
16:    $\mathit{ncss}[k]\leftarrow\mathit{updates}_{arMat}(i,\mathit{ncss}[k],\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Update ancestry matrix, Alg. 1
17:   end for
18:  end for
19:end for
20:return $\mathit{ncss}[K]$
A.3 Experiments for $K>3$ decision trees
The experimental results for $K>3$ decision trees are given in Table 7.